1a) The time needed for college students to complete a certain paper and pencil maze follows a normal distribution with a mean of 30 seconds and a standard deviation of 3 seconds. You wish to see if the mean time ? is changed by vigorous exercise, so you have a group of nine college students exercise vigorously for 20 minutes and then complete the maze.

The Test Statistic and its Distribution

In example (1a) above, it took the college students an average of x = 31.2 seconds to complete the maze. What is the distribution of the sample mean for this example? Find the corresponding test statistic and its distribution.
Answer #1

Ans:

Sampling distribution of sample means will follow normal distribution with mean=30 and standard deviation=3/sqrt(9)=1

Test statistic:

z=(31.2-30)/(3/sqrt(9))

z=1.20

p-value=2*P(z>1.20)=0.230